# Jacobian Descent For Multi-Objective Optimization

### v1

@@ 7 Conclusion
In this paper, we have defined the JD algorithm, characterized by an aggregator. Moreover, we have formulated desired aggregator properties and proposed $\mathcal{A}_{\text{UPGrad}}$ to alleviate the weaknesses of existing methods. Then, we have presented several possible applications for JD and aggregators. In particular, we have introduced IWRM, a novel learning paradigm minimizing the vector of per-instance losses. An empirical assessment of IWRM on some usual image classification datasets shows promising results and hints that it could be used as a general-purpose learning paradigm. As speed is the main limitation of JD, we have given a path towards an efficient implementation. We think that a fast algorithm to compute the Gramian of the Jacobian could unlock the full potential of JD. An important takeaway from our experiments is that considering an optimization problem in its entire dimensionality can be beneficial to its resolution. We thus hope that our work enables more multi-objective optimization problems to be considered as such.
Limitations and future directions
Our experimentation has some important limitations. First, we only evaluate JD on the minimization of the IWRM objective. It would be essential to develop proper benchmarks to compare aggregators on a wide variety of problems. Ideally, such problems should involve substantially conflicting objectives, e.g. multi-task learning with inherently competing or even adversarial tasks. Then, we have limited our scope to the comparison of optimization speeds, disregarding generalization. While this simplifies the experiments and makes the comparison precise and rigorous, optimization and generalization are sometimes intertwined. We thus believe that future works should focus on both aspects.
@@ Acknowledgments
We would like to express our sincere thanks to Scott Pesme, Emre Telatar, Matthieu Buot de l’Épine, Adrien Vandenbroucque, Alix Jeannerot and Damian Dudzicz for their careful and thorough review. The many insightful discussions that we shared with them were essential to this project.
@@ A.1 Supplementary theoretical results
Recall that a function ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to\mathbb{R}^{m}$ is -convex if for all ${\boldsymbol{x}},{\boldsymbol{y}}\in\mathbb{R}^{n}$ and any $\lambda\in[0,1]$,
${\boldsymbol{f}}(\lambda{\boldsymbol{x}}+(1-\lambda){\boldsymbol{y}})\preceq\lambda{\boldsymbol{f}}({\boldsymbol{x}})+(1-\lambda){\boldsymbol{f}}({\boldsymbol{y}}).$
@@
If ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is a continuously differentiable -convex function, then for any pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$, $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})({\boldsymbol{y}}-{\boldsymbol{x}})\preceq{\boldsymbol{f}}({\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}})$.
@@
$\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})({\boldsymbol{y}}-{\boldsymbol{x}})$ $=\lim_{\lambda\to 0^{+}}\frac{{\boldsymbol{f}}({\boldsymbol{x}}+\lambda({\boldsymbol{y}}-{\boldsymbol{x}}))-{\boldsymbol{f}}({\boldsymbol{x}})}{\lambda}$ $\left(\begin{subarray}{c}\text{differentiation}\end{subarray}\right)$
$\preceq\lim_{\lambda\to 0^{+}}\frac{{\boldsymbol{f}}({\boldsymbol{x}})+\lambda({\boldsymbol{f}}({\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}}))-{\boldsymbol{f}}({\boldsymbol{x}})}{\lambda}$ (-$\left(\begin{subarray}{c}\preceq\text{-convexity}\end{subarray}\right)$
$={\boldsymbol{f}}({\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}}),$
@@
Let $J\in\mathbb{R}^{m\times n}$, let ${\boldsymbol{u}}\in\mathbb{R}^{m}$ and let ${\boldsymbol{x}}\in\mathbb{R}^{n}$, then
${\boldsymbol{u}}^{\top}J{\boldsymbol{x}}\varleq\|{\boldsymbol{u}}\|\cdot\|J\|_{\text{F}}\cdot\|{\boldsymbol{x}}\|$
@@
$({\boldsymbol{u}}^{\top}J{\boldsymbol{x}})^{2}$ $\varleq\|{\boldsymbol{u}}\|^{2}\cdot\|J{\boldsymbol{x}}\|^{2}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
$=\|{\boldsymbol{u}}\|^{2}\cdot\sum_{i=1}^{m}(J_{i}^{\top}{\boldsymbol{x}})^{2}$
$\varleq\|{\boldsymbol{u}}\|^{2}\cdot\sum_{i=1}^{m}\|J_{i}\|^{2}\cdot\|{\boldsymbol{x}}\|^{2}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
$=\|{\boldsymbol{u}}\|^{2}\cdot\|J\|_{\text{F}}^{2}\cdot\|{\boldsymbol{x}}\|^{2},$
@@
Recall that a function ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to\mathbb{R}^{m}$ is $\beta$-smooth if for all ${\boldsymbol{x}},{\boldsymbol{y}}\in\mathbb{R}^{n}$,
$\|\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})\|_{\text{F}}\varleq\beta\|{\boldsymbol{x}}-{\boldsymbol{y}}\|$ (13)
@@
Let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be $\beta$-smooth, then for any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ and any ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$,
$)\displaystyle{\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}})-{\boldsymbol{f}}({\boldsymbol{y}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})({\boldsymbol{x}}-{\boldsymbol{y}})$$)\varleq\frac{\beta}{2}\|{\boldsymbol{w}}\|\cdot\|{\boldsymbol{x}}-{\boldsymbol{y}}\|^{2}$ (14)
@@
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}})-{\boldsymbol{f}}({\boldsymbol{y}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{y}})({\boldsymbol{x}}-{\boldsymbol{y}}))$
@@ A.2 Proposition 1
Let $J\in\mathbb{R}^{m\times n}$. For any ${\boldsymbol{u}}\in{\mathbb{R}}^{m}$, $\pi_{J}(J^{\top}{\boldsymbol{u}})=J^{\top}{\boldsymbol{w}}$ with
${\boldsymbol{w}}$ $\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{v}}\end{subarray}}{\boldsymbol{v}}^{\top}JJ^{\top}{\boldsymbol{v}}.$ (4)
@@
Let $J\in\mathbb{R}^{m\times n}$, $G=JJ^{\top}$, ${\boldsymbol{u}}\in{\mathbb{R}}^{m}$. For any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ satisfying
${\boldsymbol{u}}\preceq{\boldsymbol{w}}$ (15a)
${\boldsymbol{0}}\preceq G{\boldsymbol{w}}$ (15b)
${\boldsymbol{u}}^{\top}G{\boldsymbol{w}}={\boldsymbol{w}}^{\top}G{\boldsymbol{w}}$ (15c)
we have $\pi_{J}(J^{\top}{\boldsymbol{u}})=J^{\top}{\boldsymbol{w}}$. Such a ${\boldsymbol{w}}$ is the solution to
${\boldsymbol{w}}$ $\in\operatorname*{arg\,min}_{{\boldsymbol{u}}\preceq{\boldsymbol{v}}}{\boldsymbol{v}}^{\top}G{\boldsymbol{v}}.$
@@
Note that
$\pi_{J}(J^{\top}{\boldsymbol{u}})=\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{x}}\in{\mathbb{R}}^{n}:\\ {\boldsymbol{0}}\preceq J{\boldsymbol{x}}\end{subarray}}\frac{1}{2}\|{\boldsymbol{x}}-J^{\top}{\boldsymbol{u}}\|^{2}$
is a convex program. Consequently, the KKT conditions are both necessary and sufficient. The Lagragian is given by $\mathcal{L}({\boldsymbol{x}},{\boldsymbol{v}})=\frac{1}{2}\|{\boldsymbol{x}}-J^{\top}{\boldsymbol{u}}\|^{2}-{\boldsymbol{v}}^{\top}J{\boldsymbol{x}}$. The KKT conditions are then given by
$\begin{cases}\nabla_{\boldsymbol{x}}\mathcal{L}({\boldsymbol{x}},{\boldsymbol{v}})={\boldsymbol{0}}\\ {\boldsymbol{0}}\preceq{\boldsymbol{v}}\\ {\boldsymbol{0}}\preceq J{\boldsymbol{x}}\\ 0={\boldsymbol{v}}^{\top}J{\boldsymbol{x}}\end{cases}$
@@
The simple change of variable ${\boldsymbol{w}}={\boldsymbol{u}}+{\boldsymbol{v}}$ finishes the proof of the first part.
Since ${\boldsymbol{x}}=J^{\top}({\boldsymbol{u}}+{\boldsymbol{v}})$, the Wolfe dual program of $\pi_{J}(J^{\top}{\boldsymbol{u}})$ gives
${\boldsymbol{w}}$ $\in{\boldsymbol{u}}+\operatorname*{arg\,max}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{0}}\preceq{\boldsymbol{v}}}\mathcal{L}(J^{\top}({\boldsymbol{u}}+{\boldsymbol{v}}),{\boldsymbol{v}})$
$={\boldsymbol{u}}+\operatorname*{arg\,max}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{0}}\preceq{\boldsymbol{v}}}\frac{1}{2}\|J^{\top}{\boldsymbol{v}}\|^{2}-{\boldsymbol{v}}^{\top}JJ^{\top}({\boldsymbol{u}}+{\boldsymbol{v}})$
$={\boldsymbol{u}}+\operatorname*{arg\,max}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{0}}\preceq{\boldsymbol{v}}}-\frac{1}{2}{\boldsymbol{v}}^{\top}G{\boldsymbol{v}}-{\boldsymbol{v}}^{\top}G{\boldsymbol{u}}$
$={\boldsymbol{u}}+\operatorname*{arg\,min}_{{\boldsymbol{v}}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{u}}+{\boldsymbol{v}}}\frac{1}{2}({\boldsymbol{u}}+{\boldsymbol{v}})^{\top}G({\boldsymbol{u}}+{\boldsymbol{v}})$
$=\operatorname*{arg\,min}_{{\boldsymbol{v}}^{\prime}\in{\mathbb{R}}^{m}:~{}{\boldsymbol{u}}\preceq{\boldsymbol{v}}^{\prime}}\frac{1}{2}{\boldsymbol{v}}^{\prime\top}G{\boldsymbol{v}}^{\prime},$
@@ A.3 Theorem 1
Let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be a $\beta$-smooth and -convex function. Suppose that the Pareto front ${\boldsymbol{f}}(X^{\ast})$ is bounded and that for any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, there is ${\boldsymbol{x}}^{\ast}\in X^{\ast}$ such that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}({\boldsymbol{x}})$. Let ${\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, and for all $t\in\mathbb{N}$, ${\boldsymbol{x}}_{t+1}={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$, with $\eta=\frac{1}{\beta\sqrt{m}}$. Let ${\boldsymbol{w}}_{t}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$ as per (5), i.e. $\mathcal{A}_{\text{UPGrad}}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t}))=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})^{\top}\cdot{\boldsymbol{w}}_{t}$. If ${\boldsymbol{w}}_{t}$ is bounded, then ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})$ for some ${\boldsymbol{x}}^{\ast}\in X^{\ast}$. In other words, ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to the Pareto front.
@@
Let $J\in\mathbb{R}^{m\times n}$ and ${\boldsymbol{w}}=\frac{1}{m}\sum_{i=1}^{m}{\boldsymbol{w}}_{i}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(J)$ as per (5). Let, as usual, $G=JJ^{\top}$, then,
${\boldsymbol{w}}^{\top}G{\boldsymbol{w}}\varleq{\boldsymbol{1}}^{\top}G{\boldsymbol{w}}.$
@@
$({\boldsymbol{u}}^{\top}G{\boldsymbol{v}})^{2}$ $=\langle{\boldsymbol{u}},{\boldsymbol{v}}\rangle^{2}$
$\varleq\langle{\boldsymbol{u}},{\boldsymbol{u}}\rangle\cdot\langle{\boldsymbol{v}},{\boldsymbol{v}}\rangle$
$={\boldsymbol{u}}^{\top}G{\boldsymbol{u}}\cdot{\boldsymbol{v}}^{\top}G{\boldsymbol{v}}.$
@@
$\frac{1}{m^{2}}\sum_{i,j}\sqrt{{\boldsymbol{w}}_{i}^{\top}G{\boldsymbol{w}}_{i}}\cdot\sqrt{{\boldsymbol{w}}_{j}^{\top}G{\boldsymbol{w}}_{j}}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
@@
$~{}\sum_{i}\frac{1}{m}\left(\sqrt{{\boldsymbol{w}}_{i}^{\top}G{\boldsymbol{w}}_{i}}\right)^{2}$ (Jensen’s $\left(\begin{subarray}{c}\text{Jensen's inequality}\end{subarray}\right)$
@@
Under the assumptions of Theorem 1, for any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$,
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))\varleq\frac{\|{\boldsymbol{w}}\|}{\beta\sqrt{m}}\left(\frac{{\boldsymbol{1}}}{2\sqrt{m}}-\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}\right)^{\top}G_{t}{\boldsymbol{w}}_{t}.$
@@
For all $t\vargeq 0$, let $J_{t}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})$, $G_{t}=J_{t}J_{t}^{\top}$. Then ${\boldsymbol{x}}_{t+1}={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(J_{t})={\boldsymbol{x}}_{t}-\eta J_{t}^{\top}{\boldsymbol{w}}_{t}$. Therefore
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
@@
Under the assumptions of Theorem 1, if ${\boldsymbol{x}}^{\ast}\in X^{\ast}$ is such that ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\varleq{\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{t})$ for all $t$, then
$\frac{1}{T}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))\varleq\frac{1}{T}\bigg{(}{\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{0})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))+\frac{\beta\sqrt{m}}{2}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}\bigg{)}.$ (16)

### v2

@@ 7 Conclusion
In this paper, we introduced Jacobian descent (JD), a multi-objective optimization algorithm defined by some aggregator that maps the Jacobian to an update direction. We identified desirable properties for aggregators and proposed $\mathcal{A}_{\text{UPGrad}}$, addressing the limitations of existing methods while providing stronger convergence guarantees. We also highlighted potential applications of JD and proposed IWRM, a novel learning paradigm considering the loss of each training example as a distinct objective. Given its promising empirical results, we believe this paradigm deserves further attention. Additionally, we see potential for $\mathcal{A}_{\text{UPGrad}}$ beyond JD, as a linear algebra tool for combining conflicting vectors in broader contexts. As speed is the primary limitation of JD, we have outlined an algorithm for efficiently computing the Gramian of the Jacobian, which could unlock JD’s full potential. We hope this work serves as a foundation for future research in multi-objective optimization and encourages a broader adoption of these methods.
Limitations and future directions.
Our experimentation has some limitations. First, we only evaluate JD on IWRM, a setting with moderately conflicting objectives. It would be essential to develop proper benchmarks to compare aggregators on a wide variety of problems. Ideally, such problems should involve substantially conflicting objectives, e.g. multi-task learning with inherently competing or even adversarial tasks. Then, we have limited our scope to the comparison of optimization speeds, disregarding generalization. While this simplifies the experiments and makes the comparison rigorous, optimization and generalization are sometimes intertwined. We thus believe that future works should focus on both aspects.
@@ Acknowledgments
We would like to express our sincere thanks to Scott Pesme, Emre Telatar, Matthieu Buot de l’Épine, Adrien Vandenbroucque, Ye Zhu, Alix Jeannerot, and Damian Dudzicz for their careful and thorough review. The many insightful discussions that we shared with them were essential to this project.
@@ A.1 Supplementary theoretical results
Recall that a function ${\bm{f}}:{\mathbb{R}}^{n}\to\mathbb{R}^{m}$ is -convex if for all ${\bm{x}},{\bm{y}}\in\mathbb{R}^{n}$ and any $\lambda\in[0,1]$,
${\bm{f}}\big{(}\lambda{\bm{x}}+(1-\lambda){\bm{y}}\big{)}\preceq\lambda{\bm{f}}({\bm{x}})+(1-\lambda){\bm{f}}({\bm{y}}).$
@@
If ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is a continuously differentiable -convex function, then for any pair of vectors ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{n}$, $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})({\bm{y}}-{\bm{x}})\preceq{\bm{f}}({\bm{y}})-{\bm{f}}({\bm{x}})$.
@@
$\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})({\bm{y}}-{\bm{x}})$ $=\lim_{\lambda\to 0^{+}}\frac{{\bm{f}}\big{(}{\bm{x}}+\lambda({\bm{y}}-{\bm{x}})\big{)}-{\bm{f}}({\bm{x}})}{\lambda}$ $\left(\begin{subarray}{c}\text{differentiation}\end{subarray}\right)$
$\preceq\lim_{\lambda\to 0^{+}}\frac{{\bm{f}}({\bm{x}})+\lambda\big{(}{\bm{f}}({\bm{y}})-{\bm{f}}({\bm{x}})\big{)}-{\bm{f}}({\bm{x}})}{\lambda}$ (-$\left(\begin{subarray}{c}\preceq\text{-convexity}\end{subarray}\right)$
$={\bm{f}}({\bm{y}})-{\bm{f}}({\bm{x}}),$
@@
Let $J\in\mathbb{R}^{m\times n}$, let ${\bm{u}}\in\mathbb{R}^{m}$ and let ${\bm{x}}\in\mathbb{R}^{n}$, then
${\bm{u}}^{\top}J{\bm{x}}\varleq\|{\bm{u}}\|\cdot\|J\|_{\text{F}}\cdot\|{\bm{x}}\|$
@@
$\left({\bm{u}}^{\top}J{\bm{x}}\right)^{2}$ $\varleq\|{\bm{u}}\|^{2}\cdot\|J{\bm{x}}\|^{2}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
$=\|{\bm{u}}\|^{2}\cdot\sum_{i\in[m]}\left(J_{i}^{\top}{\bm{x}}\right)^{2}$
$\varleq\|{\bm{u}}\|^{2}\cdot\sum_{i\in[m]}\|J_{i}\|^{2}\cdot\|{\bm{x}}\|^{2}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
$=\|{\bm{u}}\|^{2}\cdot\|J\|_{\text{F}}^{2}\cdot\|{\bm{x}}\|^{2},$
@@
Recall that a function ${\bm{f}}:{\mathbb{R}}^{n}\to\mathbb{R}^{m}$ is $\beta$-smooth if for all ${\bm{x}},{\bm{y}}\in\mathbb{R}^{n}$,
$\big{\|}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})-\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{y}})\big{\|}_{\text{F}}\varleq\beta\|{\bm{x}}-{\bm{y}}\|$ (13)
@@
Let ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be $\beta$-smooth, then for any ${\bm{w}}\in{\mathbb{R}}^{m}$ and any ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{n}$,
${\bm{w}}^{\top}\big{(}{\bm{f}}({\bm{x}})-{\bm{f}}({\bm{y}})-\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{y}})({\bm{x}}-{\bm{y}})\big{)}\varleq\frac{\beta}{2}\|{\bm{w}}\|\cdot\|{\bm{x}}-{\bm{y}}\|^{2}$ (14)
@@
${\bm{w}}^{\top}\big{(}{\bm{f}}({\bm{x}})-{\bm{f}}({\bm{y}})-\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{y}})({\bm{x}}-{\bm{y}})\big{)}$
@@ A.2 Proposition 1
Let $J\in\mathbb{R}^{m\times n}$. For any ${\bm{u}}\in{\mathbb{R}}^{m}$, $\pi_{J}(J^{\top}{\bm{u}})=J^{\top}{\bm{w}}$ with
${\bm{w}}$ $\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\bm{v}}\in{\mathbb{R}}^{m}:~{}{\bm{u}}\preceq{\bm{v}}\end{subarray}}{\bm{v}}^{\top}JJ^{\top}{\bm{v}}.$ (5)
@@
Let $J\in\mathbb{R}^{m\times n}$, $G=JJ^{\top}$, ${\bm{u}}\in{\mathbb{R}}^{m}$. For any ${\bm{w}}\in{\mathbb{R}}^{m}$ satisfying
${\bm{u}}\preceq{\bm{w}}$ (15a)
${\bm{0}}\preceq G{\bm{w}}$ (15b)
${\bm{u}}^{\top}G{\bm{w}}={\bm{w}}^{\top}G{\bm{w}}$ (15c)
we have $\pi_{J}(J^{\top}{\bm{u}})=J^{\top}{\bm{w}}$. Such a ${\bm{w}}$ is the solution to
${\bm{w}}$ $\in\operatorname*{arg\,min}_{{\bm{u}}\preceq{\bm{v}}}{\bm{v}}^{\top}G{\bm{v}}.$
@@
The projection
$\pi_{J}(J^{\top}{\bm{u}})=\operatorname*{arg\,min}_{\begin{subarray}{c}{\bm{x}}\in{\mathbb{R}}^{n}:\\ {\bm{0}}\preceq J{\bm{x}}\end{subarray}}\frac{1}{2}\|{\bm{x}}-J^{\top}{\bm{u}}\|^{2}$
is a convex program. Consequently, the KKT conditions are both necessary and sufficient. The Lagragian is given by $\mathcal{L}({\bm{x}},{\bm{v}})=\frac{1}{2}\|{\bm{x}}-J^{\top}{\bm{u}}\|^{2}-{\bm{v}}^{\top}J{\bm{x}}$. The KKT conditions are then given by
$\begin{cases}\nabla_{\bm{x}}\mathcal{L}({\bm{x}},{\bm{v}})={\bm{0}}\\ {\bm{0}}\preceq{\bm{v}}\\ {\bm{0}}\preceq J{\bm{x}}\\ 0={\bm{v}}^{\top}J{\bm{x}}\end{cases}$
@@
The simple change of variable ${\bm{w}}={\bm{u}}+{\bm{v}}$ finishes the proof of the first part.
Since ${\bm{x}}=J^{\top}({\bm{u}}+{\bm{v}})$, the Wolfe dual program of $\pi_{J}(J^{\top}{\bm{u}})$ gives
${\bm{w}}$ $\in{\bm{u}}+\operatorname*{arg\,max}_{{\bm{v}}\in{\mathbb{R}}^{m}:~{}{\bm{0}}\preceq{\bm{v}}}\mathcal{L}\big{(}J^{\top}({\bm{u}}+{\bm{v}}),{\bm{v}}\big{)}$
$={\bm{u}}+\operatorname*{arg\,max}_{{\bm{v}}\in{\mathbb{R}}^{m}:~{}{\bm{0}}\preceq{\bm{v}}}\frac{1}{2}\left\|J^{\top}{\bm{v}}\right\|^{2}-{\bm{v}}^{\top}JJ^{\top}({\bm{u}}+{\bm{v}})$
$={\bm{u}}+\operatorname*{arg\,max}_{{\bm{v}}\in{\mathbb{R}}^{m}:~{}{\bm{0}}\preceq{\bm{v}}}-\frac{1}{2}{\bm{v}}^{\top}G{\bm{v}}-{\bm{v}}^{\top}G{\bm{u}}$
$={\bm{u}}+\operatorname*{arg\,min}_{{\bm{v}}\in{\mathbb{R}}^{m}:~{}{\bm{u}}\preceq{\bm{u}}+{\bm{v}}}\frac{1}{2}({\bm{u}}+{\bm{v}})^{\top}G({\bm{u}}+{\bm{v}})$
$=\operatorname*{arg\,min}_{{\bm{v}}^{\prime}\in{\mathbb{R}}^{m}:~{}{\bm{u}}\preceq{\bm{v}}^{\prime}}\frac{1}{2}{\bm{v}}^{\prime\top}G{\bm{v}}^{\prime},$
@@ A.3 Theorem 1
Let ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ be a $\beta$-smooth and -convex function. Suppose that the Pareto front ${\bm{f}}(X^{\ast})$ is bounded and that for any ${\bm{x}}\in{\mathbb{R}}^{n}$, there is ${\bm{x}}^{\ast}\in X^{\ast}$ satisfying ${\bm{f}}({\bm{x}}^{\ast})\preceq{\bm{f}}({\bm{x}})$. Let ${\bm{x}}_{1}\in{\mathbb{R}}^{n}$, and for all $t\in\mathbb{N}$, ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})\big{)}$, with $\eta=\frac{1}{\beta\sqrt{m}}$. Let ${\bm{w}}_{t}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})\big{)}$ as per (6), i.e. $\mathcal{A}_{\text{UPGrad}}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})\big{)}=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})^{\top}\cdot{\bm{w}}_{t}$. If ${\bm{w}}_{t}$ is bounded, then ${\bm{f}}({\bm{x}}_{t})$ converges to ${\bm{f}}({\bm{x}}^{\ast})$ for some ${\bm{x}}^{\ast}\in X^{\ast}$. In other words, ${\bm{f}}({\bm{x}}_{t})$ converges to the Pareto front.
@@
Let $J\in\mathbb{R}^{m\times n}$ and ${\bm{w}}=\frac{1}{m}\sum_{i=1}^{m}{\bm{w}}_{i}$ be the weights defining $\mathcal{A}_{\text{UPGrad}}(J)$ as per (6). Let, as usual, $G=JJ^{\top}$, then,
${\bm{w}}^{\top}G{\bm{w}}\varleq{\bm{1}}^{\top}G{\bm{w}}.$
@@
$({\bm{u}}^{\top}G{\bm{v}})^{2}$ $=\langle{\bm{u}},{\bm{v}}\rangle^{2}$
$\varleq\langle{\bm{u}},{\bm{u}}\rangle\cdot\langle{\bm{v}},{\bm{v}}\rangle$
$={\bm{u}}^{\top}G{\bm{u}}\cdot{\bm{v}}^{\top}G{\bm{v}}.$
@@
$\frac{1}{m^{2}}\sum_{i,j}\sqrt{{\bm{w}}_{i}^{\top}G{\bm{w}}_{i}}\cdot\sqrt{{\bm{w}}_{j}^{\top}G{\bm{w}}_{j}}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
@@
$~{}\sum_{i}\frac{1}{m}\left(\sqrt{{\bm{w}}_{i}^{\top}G{\bm{w}}_{i}}\right)^{2}$ (Jensen’s $\left(\begin{subarray}{c}\text{Jensen's inequality}\end{subarray}\right)$
@@
Under the assumptions of Theorem 1, for any ${\bm{w}}\in{\mathbb{R}}^{m}$ and any $t\in\mathbb{N}$,
${\bm{w}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t+1})-{\bm{f}}({\bm{x}}_{t})\big{)}\varleq\frac{\|{\bm{w}}\|}{\beta\sqrt{m}}\left(\frac{{\bm{1}}}{2\sqrt{m}}-\frac{{\bm{w}}}{\|{\bm{w}}\|}\right)^{\top}G_{t}{\bm{w}}_{t}.$
@@
For all $t\in\mathbb{N}$, let $J_{t}=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})$, $G_{t}=J_{t}J_{t}^{\top}$. Then ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(J_{t})={\bm{x}}_{t}-\eta J_{t}^{\top}{\bm{w}}_{t}$. Therefore
${\bm{w}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t+1})-{\bm{f}}({\bm{x}}_{t})\big{)}$
@@
Under the assumptions of Theorem 1, if ${\bm{x}}^{\ast}\in X^{\ast}$ satisfies ${\bm{1}}^{\top}{\bm{f}}({\bm{x}}^{\ast})\varleq{\bm{1}}^{\top}{\bm{f}}({\bm{x}}_{t})$ for all $t\in\mathbb{N}$, then
$\frac{1}{T}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}^{\ast})\big{)}\varleq\frac{1}{T}\bigg{(}{\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast})\big{)}+\frac{\beta\sqrt{m}}{2}\left\|{\bm{x}}_{1}-{\bm{x}}^{\ast}\right\|^{2}\bigg{)}.$ (16)